# One Model for All Tasks: Leveraging Efficient World Models in Multi-Task Planning

### v1

@@ D.1 Benchmark Setup
Figure 12: First five steps of expert interaction trajectories for Zork1 and Detective. At each step, the environment supplies a set of valid candidate actions (used in our experiments); however, additional legal actions outside this set may also be executed, so the action space is, in principle, unbounded.
@@
As shown in Figure 13, the multitask ScaleZero attains returns comparable to the single-task UniZero baseline on most Jericho games, with learning curves exhibiting an “easy-first, hard-later” progression. Specifically, ScaleZero converges faster on Acorncourt, Omniquest, and Detective; on the most challenging Zork1, it initially lags but accelerates after  150k environment interactions to reach parity with UniZero. We hypothesize that this stems from multi-task training inducing transferable language priors and interaction routines, which help filter feasible commands, compress the open-ended action space, and reduce fruitless exploration.
@@ E.1 Experimental Analysis
Given the strong performance of Mixture-of-Experts (MoE) in multitask settings, a natural question arises: Why MoE? We aim to investigate, both experimentally and theoretically, the underlying mechanisms behind MoE’s superior performance in multitask reinforcement learning. Specifically, our experimental setup is as follows.
@@ E.1.1 Experiment 1: Analyzing Gradient Conflicts in MoE-based Transformers
(1). Naive Transformer: The backbone consists of four standard Transformer blocks.
(2). MoE-based Transformer: The backbone also consists of four Transformer blocks, but the MLP layer in each block is replaced with an MoE layer, which comprises one shared expert and eight non-shared experts(Liu et al., 2024b). Shared experts provide cross-task general representations, enhancing generalization ability and ensuring training stability, while non-shared experts learn task-specific representations to strengthen the model’s discriminability and task adaptability. During the forward pass, all non-shared experts are selectively activated by a sparse gating network, which determines which specific expert to use for each input.
@@
where $\nabla\theta_{i}$ and $\nabla\theta_{j}$ denote the gradients of the $i$-th and $j$-th tasks, respectively. A higher value of the Max Gradient Conflict represents a greater degree of gradient conflict. We choose the maximum pairwise cosine similarity because it directly identifies the most severe gradient conflict between any two tasks. Unlike averaging operation, which can hide critical issues, the maximum value pinpoints the ’bottleneck pair’ that most significantly impedes stable multi-task training and overall convergence, even if other tasks cooperate well.
@@
Observation 1: As shown in Figure˜14(a), the MoE-based Transformer exhibits fewer gradient conflicts than the MLP counterpart. (b) further shows that introducing MoE reduces gradient conflicts at the layer input, implying that MoE alleviates conflicts in other components to some extent. In the encoder, however, the conflict levels of MoE and MLP remain comparable. We argue that the main reason is that the encoder extracts general representations, which inherently exhibit fewer gradient conflicts.
@@ E.1.2 Experiment 2: Investigating MoE Gating Mechanisms
We are particularly interested in the internal gating mechanism of MoE. Previous studiesChen et al. have shown that, under supervised learning, MoE can implicitly uncover latent cluster structures within input space. However, when faced with non-stationary data distributions generated through agent-environment interactions, it remains unclear whether MoE experts still differentiate effectively. To answer the question, we analyze the entropy of expert selection distributions, which quantifies the uncertainty of the choice of which expert to use for a task, with low entropy indicating high specialization and decisive selection, while high entropy implies uncertainty or an average utilization across multiple experts. We also recorded the Wasserstein distance Rüschendorf (1985) between the expert selection distributions of different tasks, where smaller values indicate a greater proximity to the expert selection of two tasks. We aim to quantify this relationship to find the underlying connections in expert selection among various tasks. The specific experimental steps are as follows:
@@
We consider two sizes $S$: immediate = 100, short = 1,000.555Note that at the very beginning of training, when the window cannot be fully populated, we use the available data within the window, which may lead to larger fluctuations at early steps. We present the entropy results in Figure˜15. Due to space limitations, we show in Figure˜18 the Wasserstein distances between expert selections of different tasks at multiple training stages.
Figure 15: Line plot showing the evolution of expert selection entropy in a multi-task learning setting with eight tasks. The dashed lines correspond to the entropy values of individual tasks, and the solid red line represents the aggregated entropy across all tasks. Higher entropy reflects more uniform and uncertain expert utilization, while lower entropy reflects more concentrated and specialized expert selection.
Observation 2: MoE differentiation and expert specialization. As shown in Figure˜15, as tasks progress, the entropy of the expert distribution gradually decreases, indicating a reduction in distributional uncertainty. This suggests that expert selection becomes concentrated on a few outcomes, reflecting lower uncertainty. Furthermore, in Figure˜18, we visualize the expert selection distributions of different tasks at multiple training stages, offering additional evidence of the progressive specialization in expert utilization.
@@ E.1.3 Experiment 3: Analyzing Gradient Conflicts between shared expert and non-shared expert
Figure 16: Gradient conflicts across experts in MoE training. The plot shows the evolution of gradient conflicts for the shared expert and 8 non-shared experts during training interations. The logarithmic y-axis reveals that different experts experience varying levels of gradient conflicts, with the shared expert and individual experts (expert0–expert7) demonstrating distinct conflict patterns, indicating effective expert specialization in the multi-task learning framework.
Observation 3: As shown in Figure˜16, the shared expert exhibits significantly higher gradient conflicts compared to the task-specific experts. Among the non-shared experts, the level of conflicts does not show substantial differences. Overall, the shared expert accumulates even more conflicts than the entire MoE layer, indicating that most gradient conflicts within the MoE layer are concentrated on the shared expert. The introduction of several task-specific experts effectively reduces the overall gradient conflicts of the MoE layer, which is consistent with our theoretical analysis in Corollary˜E.4.
We further investigated the conflicts among different experts within MoE. Results is shown in Figure˜16 We found that shared experts bear most of the gradient conflicts, whereas individual (non-shared) experts experience almost no conflict. A plausible explanation is that in a standard gating MoE, different samples are routed to different experts, so each expert primarily receives gradient updates from a specific type of sample. This effectively performs an implicit task partitioning, leading to generally consistent gradient directions within the same expert and minimal conflicts. In contrast, in shared-expert MoE, all samples pass through the shared experts, whose parameters must adapt simultaneously to multiple tasks, diverse semantics, and even different domain distributions. This can result in highly inconsistent gradient directions, significantly increasing conflicts and making it difficult for shared experts to balance diverse task requirements during optimization.This observation naturally raises a simple question: is the alleviation of gradient conflicts mainly due to the sparse selection by the gating network? We provide a theoretical justification from the perspective of gating.
@@ E.2 Theoretical Analysis
Therefore, the key to mitigating gradient conflict lies in the router’s ability to foster specialization among experts, ensuring that tasks with potentially conflicting objectives are handled by distinct, differentiated experts.
This naturally raises a critical question: Can a Mixture-of-Experts architecture indeed achieve this, and if so, what are the underlying mechanisms that enable this capability? In the following, we answer this question affirmatively. By building upon existing foundational theories of Mixture-of-ExpertsChen et al. , we present a formal derivation that explains precisely when and how this is achieved.
@@ E.2.1 Analysis
Previews workChen et al. theoretically demonstrated that, in a single-task supervised learning setting, when the input distribution exhibits distinguishable cluster structures, the sparse gating Mixture-of-Experts (MoE) router can automatically learn the cluster center features and route samples to the most suitable expert, thereby significantly improving performance. We further hypothesize that, in multi-task learning, the input spaces of different tasks naturally correspond to distinct clusters and the signal for the cluster center can be task ID or the latent pattern on the state space. Under this assumption, the conclusions from the original work can be directly extended to the multi-task setting.
Next, we adapt Chen et al. theoretical analysis to the reinforcement learning context using a tabular example. Specifically, we treat each task as an independent Markov Decision Process (MDP) and construct a multi-task MDP set with a shared state space but task-specific transitions. In this setting, we derive the gradient conflicts that arise in the MoE layer.
@@
E.2.2 Problem setting: multi-task MDP with task-specific latent clusters
@@ E.2.3 Learning Dynamics Analysis
We follow an analysis path similar to prior workChen et al. , dividing the learning process into two stages: an early expert exploration stage and a later router learning stage.
@@
According to [Chen et al. , Lemma E.3], the zero initialization of the router implies that at the beginning of training ($t=0$),
@@
Under gradient descent updates, the weight vector $w_{m}$ of expert $m$ will predominantly grow along the direction of its preferred basis vector $v_{k_{m}^{*}}$.[Chen et al. , Lemma E.5] In other words, during training, an expert progressively specializes in modeling a specific cluster $c_{k}$. Building upon this, we derive an upper bound on the gradient conflicts during the Expert Exploration Stage.
@@ 2. For the full-layer MoE with sparse router, the expected gradient conflict is upper bounded by
During the router-learning stage, the sparse MoE layer exhibits two decisive properties that justify its use in a multi-task setting.
@@
iterations, its weight vectors $\bm{\theta}_{m}$ become strongly aligned with the cluster-center signal $\mathbf{c}_{k}$ of every task $k$ , while simultaneously suppressing spurious correlations with label signals and noise.[Chen et al. -lemma E.14] Consequently, any input drawn from task $k$ is routed to the corresponding subset of experts $\mathcal{M}_{k}$ with probability $1-o(\frac{1}{d})$, even though the model is never given explicit task labels. [Chen et al. -Lemma E.18]
Second, each expert in $\mathcal{M}_{k}$ remains tightly specialised to its assigned task, but only chance-level performance on all other tasks. [Chen et al. -Lemma E.12,Lemma 5.2].
Taken together, these two properties guarantee automatic task separation and expert specialisation without external supervision, making the MoE layer an effective backbone for multi-task problems in which tasks form separable clusters in the input space.
@@
E.2.4 Task Embedding for Improved Task Separation in Multi-Task Learning
In multi-task learning, tasks often share overlapping input spaces and underlying features (e.g., sentiment classification tasks using the same text encoder), making it difficult for the routing network to distinguish tasks. This overlap leads to routing confusion, where data from one task may be misrouted to experts specialized in other tasks, reducing performance. To address this, task embedding introduces a learnable vector $e_{\tau}$ for each task, which can be combined with input features. Since $e_{\tau}$ is orthogonal across tasks (or constrained via a regularization term), it helps the routing network more clearly distinguish between tasks, even when inputs are similar. Task embedding also acts as a regularization term, reducing expert load imbalance and improving the stability and performance of multi-task MoE training.
Our theoretical analysis suggests that introducing distinct task embeddings provides crucial clustering signals, enabling the router to more effectively differentiate between tasks. This, in turn, mitigates gradient conflicts and enhances overall performance. However, our empirical results, as presented in Figure˜3, show that the inclusion of naive task embeddings did not yield a significant improvement. We hypothesize that this is due to the absence of explicit constraints to maintain separation between the embeddings during optimization, which could lead to their collapse in the later stages of training. For future work, we plan to investigate more effective methods for explicitly injecting task-specific information. The goal is to empower the router to autonomously leverage the unique priors of each task, thereby further advancing multi-task learning performance.
@@ E.2.4 Task Embedding for Improved Task Separation in Multi-Task Learning
Figure 19: T-SNE Visualization. Each point denotes a 2D projection of a 768-dimensional latent embedding extracted from the world model. Points are colored by environment ID, with different colors representing different Atari games. The visualization reveals the learned representational structure: similar game states cluster together in the latent space, while distinct environments form separate groups. This indicates that the model captures task-specific features while preserving meaningful separation across different games. The figure contains 8,000 data points collected from 8 distributed training processes, where each point corresponds to a single time step from the observation sequence (400 samples × 20 time steps per sample).

### v2

@@ D.1 Benchmark Setup
(c) Omniquest
(d) Acorncourt
Figure 14: First five steps of expert interaction trajectories across environments. At each step, the environment supplies a set of valid candidate actions (used in our experiments); however, additional legal actions outside this set may also be executed, so the action space is, in principle, unbounded.
@@
As shown in Figure 15, the multitask ScaleZero attains returns comparable to the single-task UniZero baseline on most Jericho games, with learning curves exhibiting an “easy-first, hard-later” progression. Specifically, ScaleZero converges faster on Acorncourt, Omniquest, and Detective; on the most challenging Zork1, it initially lags but accelerates after  150k environment interactions to reach parity with UniZero. We hypothesize that this stems from multitask training inducing transferable language priors and interaction routines, which help filter feasible commands, compress the open-ended action space, and reduce fruitless exploration.
@@ E.1 Experimental Analysis
To understand why Mixture-of-Experts (MoE) architectures excel in multitask settings, this study investigates the core mechanisms driving their performance. Focusing on multitask reinforcement learning, we conduct a dual analysis combining theoretical inquiry with empirical validation. The following section details the experimental protocol designed for this purpose.
@@ E.1.1 Experiment 1: Analyzing Gradient Conflicts in MoE-based Transformers
(1) Naive Transformer: The backbone consists of four standard Transformer blocks.
(2) MoE-based Transformer: The backbone also consists of four Transformer blocks, but the MLP layer in each block is replaced with an MoE layer, which comprises one shared expert and eight non-shared experts(Liu et al., 2024). Shared experts provide cross-task general representations, enhancing generalization ability and ensuring training stability, while non-shared experts learn task-specific representations to strengthen the model’s discriminability and task adaptability. During the forward pass, all non-shared experts are selectively activated by a sparse gating network, which determines which specific expert to use for each input.
@@
where $\nabla\theta_{i}$ and $\nabla\theta_{j}$ denote the gradients of the $i$-th and $j$-th tasks, respectively. A higher value of the Max Gradient Conflict represents a greater degree of gradient conflict. We choose the maximum pairwise cosine similarity because it directly identifies the most severe gradient conflict between any two tasks. Unlike averaging operation, which can hide critical issues, the maximum value pinpoints the ’bottleneck pair’ that most significantly impedes stable multitask training and overall convergence, even if other tasks cooperate well.
@@
Observation 1: As shown in Figure˜16(a), the MoE-based Transformer exhibits fewer gradient conflicts than the MLP counterpart. (b) further shows that introducing MoE reduces gradient conflicts at the layer input, implying that MoE alleviates conflicts in other components to some extent. In the encoder, however, the conflict levels of MoE and MLP remain comparable. We argue that the main reason is that the encoder extracts general representations, which inherently exhibit fewer gradient conflicts.
@@ E.1.2 Experiment 2: Investigating MoE Gating Mechanisms
We are particularly interested in the internal gating mechanism of MoE. Previous studiesChen et al. (2022) have shown that, under supervised learning, MoE can implicitly uncover latent cluster structures within input space. However, when faced with non-stationary data distributions generated through agent-environment interactions, it remains unclear whether MoE experts still differentiate effectively. To answer the question, we analyze the entropy of expert selection distributions, which quantifies the uncertainty of the choice of which expert to use for a task, with low entropy indicating high specialization and decisive selection, while high entropy implies uncertainty or an average utilization across experts. We also record the Wasserstein distance  (Rüschendorf, 1985) between the expert selection distributions of different tasks, where smaller values indicate a greater proximity to the expert selection of two tasks. We aim to quantify this relationship to find the underlying connections in expert selection among various tasks. The specific experimental steps are as follows:
@@
We consider two sizes $S$: immediate = 100, short = 1,000.444Note that at the very beginning of training, when the window cannot be fully populated, we use the available data within the window, which may lead to larger fluctuations at early steps. We present the entropy results in Figure˜17. Due to space limitations, we show in Figure˜20 the Wasserstein distances between expert selections of different tasks at multiple training stages.
Figure 17: Line plot showing the evolution of expert selection entropy in a multitask learning setting with eight tasks. The dashed lines correspond to the entropy values of individual tasks, and the solid red line represents the aggregated entropy across all tasks. Higher entropy reflects more uniform and uncertain expert utilization, while lower entropy reflects more concentrated and specialized expert selection.
Observation 2: MoE differentiation and expert specialization. As shown in Figure˜17, as tasks progress, the entropy of the expert distribution gradually decreases, indicating a reduction in distributional uncertainty. This suggests that expert selection becomes concentrated on a few outcomes, reflecting lower uncertainty. Furthermore, in Figure˜20, we visualize the expert selection distributions of different tasks at multiple training stages, offering additional evidence of the progressive specialization in expert utilization.
@@ E.1.3 Experiment 3: Analyzing Gradient Conflicts between shared expert and non-shared expert
Figure 18: Gradient conflicts across experts in MoE training. The plot shows the evolution of gradient conflicts for the shared expert and 8 non-shared experts during training interations. The logarithmic y-axis reveals that different experts experience varying levels of gradient conflicts, with the shared expert and individual experts (expert0–expert7) demonstrating distinct conflict patterns, indicating effective expert specialization in the multitask learning framework.
Observation 3: As shown in Figure˜18, the shared expert exhibits significantly higher gradient conflicts compared to the task-specific experts. Among the non-shared experts, the level of conflicts does not show substantial differences. Overall, the shared expert accumulates even more conflicts than the entire MoE layer, indicating that most gradient conflicts within the MoE layer are concentrated on the shared expert. The introduction of several task-specific experts effectively reduces the overall gradient conflicts of the MoE layer, which is consistent with our theoretical analysis in Corollary˜E.4.
We further investigated the conflicts among different experts within MoE. Results is shown in Figure˜18 We found that shared experts bear most of the gradient conflicts, whereas individual (non-shared) experts experience almost no conflict. A plausible explanation is that in a standard gating MoE, different samples are routed to different experts, so each expert primarily receives gradient updates from a specific type of sample. This effectively performs an implicit task partitioning, leading to generally consistent gradient directions within the same expert and minimal conflicts. In contrast, in shared-expert MoE, all samples pass through the shared experts, whose parameters must adapt simultaneously to multiple tasks, diverse semantics, and even different domain distributions. This can result in highly inconsistent gradient directions, significantly increasing conflicts and making it difficult for shared experts to balance diverse task requirements during optimization.This observation naturally raises a simple question: is the alleviation of gradient conflicts mainly due to the sparse selection by the gating network? We provide a theoretical justification from the perspective of gating.
@@ E.2 Theoretical Analysis
Therefore, the key to mitigating gradient conflict lies in the router’s ability to foster specialization among experts, ensuring that tasks with potentially conflicting objectives are handled by distinct, differentiated experts. This naturally raises a critical question: Can a Mixture-of-Experts architecture indeed achieve this, and if so, what are the underlying mechanisms that enable this capability? In the following, we answer this question affirmatively. By building upon existing foundational theories of Mixture-of-ExpertsChen et al. (2022), we present a formal derivation that explains precisely when and how this is achieved.
@@ E.2.1 Analysis
Previews workChen et al. (2022) theoretically demonstrated that, in a single-task supervised learning setting, when the input distribution exhibits distinguishable cluster structures, the sparse gating Mixture-of-Experts (MoE) router can automatically learn the cluster center features and route samples to the most suitable expert, thereby significantly improving performance. We further hypothesize that, in multitask learning, the input spaces of different tasks naturally correspond to distinct clusters and the signal for the cluster center can be task ID or the latent pattern on the state space. Under this assumption, the conclusions from the original work can be directly extended to the multitask setting.
Next, we adapt Chen et al. (2022) theoretical analysis to the reinforcement learning context using a tabular example. Specifically, we treat each task as an independent Markov Decision Process (MDP) and construct a multitask MDP set with a shared state space but task-specific transitions. In this setting, we derive the gradient conflicts that arise in the MoE layer.
@@
E.2.2 Problem setting: multitask MDP with task-specific latent clusters
@@ E.2.3 Learning Dynamics Analysis
We follow an analysis path similar to prior workChen et al. (2022), dividing the learning process into two stages: an early expert exploration stage and a later router learning stage.
@@
According to [Chen et al. (2022), Lemma E.3], the zero initialization of the router implies that at the beginning of training ($t=0$),
@@
Under gradient descent updates, the weight vector $w_{m}$ of expert $m$ will predominantly grow along the direction of its preferred basis vector $v_{k_{m}^{*}}$.[Chen et al. (2022), Lemma E.5] In other words, during training, an expert progressively specializes in modeling a specific cluster $c_{k}$. Building upon this, we derive an upper bound on the gradient conflicts during the Expert Exploration Stage.
@@ 2. For the full-layer MoE with sparse router, the expected gradient conflict is upper bounded by
During the router-learning stage, the sparse MoE layer exhibits two decisive properties that justify its use in a multitask setting.
@@
iterations, its weight vectors $\bm{\theta}_{m}$ become strongly aligned with the cluster-center signal $\mathbf{c}_{k}$ of every task $k$ , while simultaneously suppressing spurious correlations with label signals and noise.[Chen et al. (2022)-lemma E.14] Consequently, any input drawn from task $k$ is routed to the corresponding subset of experts $\mathcal{M}_{k}$ with probability $1-o(\frac{1}{d})$, even though the model is never given explicit task labels. [Chen et al. (2022)-Lemma E.18]
Second, each expert in $\mathcal{M}_{k}$ remains tightly specialised to its assigned task, but only chance-level performance on all other tasks. [Chen et al. (2022)-Lemma E.12,Lemma 5.2].
Taken together, these two properties guarantee automatic task separation and expert specialisation without external supervision, making the MoE layer an effective backbone for multitask problems in which tasks form separable clusters in the input space.
@@
E.2.4 Task Embedding for Improved Task Separation in multitask Learning
In multitask learning, tasks often share overlapping input spaces and underlying features (e.g., sentiment classification tasks using the same text encoder), making it difficult for the routing network to distinguish tasks. This overlap leads to routing confusion, where data from one task may be misrouted to experts specialized in other tasks, reducing performance. To address this, task embedding introduces a learnable vector $e_{\tau}$ for each task, which can be combined with input features. Since $e_{\tau}$ is orthogonal across tasks (or constrained via a regularization term), it helps the routing network more clearly distinguish between tasks, even when inputs are similar. Task embedding also acts as a regularization term, reducing expert load imbalance and improving the stability and performance of multitask MoE training.
Our theoretical analysis suggests that introducing distinct task embeddings provides crucial clustering signals, enabling the router to more effectively differentiate between tasks. This, in turn, mitigates gradient conflicts and enhances overall performance. However, our empirical results, as presented in Figure˜3, show that the inclusion of naive task embeddings did not yield a significant improvement. We hypothesize that this is due to the absence of explicit constraints to maintain separation between the embeddings during optimization, which could lead to their collapse in the later stages of training. For future work, we plan to investigate more effective methods for explicitly injecting task-specific information. The goal is to empower the router to autonomously leverage the unique priors of each task, thereby further advancing multitask learning performance.
@@ E.2.4 Task Embedding for Improved Task Separation in multitask Learning
Figure 21: T-SNE Visualization of latent space. Each point denotes a 2D projection of a 768-dimensional latent embedding extracted from the world model. Points are colored by environment ID, with different colors representing different Atari games. The visualization reveals the learned representational structure: similar game states cluster together in the latent space, while distinct environments form separate groups. This indicates that the model captures task-specific features while preserving meaningful separation across different games. The figure contains 8,000 data points collected from 8 distributed training processes, where each point corresponds to a single time step from the observation sequence (400 samples × 20 time steps per sample).
Appendix F Detailed Related Work
MCTS with Learned World Models. Planning within a learned latent space, popularized by MuZero (Schrittwieser et al., 2019), has become a dominant paradigm in reinforcement learning, building upon the success of AlphaZero (Silver et al., 2016; 2017). This approach enables a model to predict its own dynamics, rewards, and policies, allowing for powerful lookahead search within a compact, learned representation (Hubert et al., 2021; Antonoglou et al., 2021; Danihelka et al., 2022; Xuan et al., 2024). Recent advancements have incorporated Transformers as the backbone for these world models (Micheli et al., 2022; Robine et al., 2023; Pu et al., 2024; Zhou et al., 2024), significantly enhancing representational capacity and the ability to model long-horizon dependencies. However, the monolithic nature of these architectures, while powerful for single tasks, becomes a critical liability in heterogeneous multitask settings. They are prone to representational interference and plasticity collapse, where a single shared model struggles to accommodate the diverse and often conflicting properties of multiple tasks. Our work confronts this fundamental architectural bottleneck by systematically investigating and proposing a more robust design.
Multi-Task Reinforcement Learning (MTRL). MTRL seeks to improve data efficiency and generalization by sharing knowledge across a distribution of tasks (Vithayathil Varghese & Mahmoud, 2020; D’Eramo et al., 2024). A common strategy is to use a shared backbone with task-specific heads to handle diverse action or observation spaces (Kumar et al., 2022; Hansen et al., 2023; Lee et al., 2022; Gallouédec et al., 2024). To mitigate inter-task interference within the shared backbone, various approaches have been proposed. Architectural methods aim to disentangle knowledge through context-based conditioning (Rakelly et al., 2019; Sodhani et al., 2021), learnable modulation modules that adapt network activations (Schmied et al., 2023), or explicit modularization with task-aware routing and parameter composition (Sun et al., 2022; He et al., 2023). In parallel, optimization-based methods focus on managing gradient conflicts at the training level by projecting gradients to avoid negative interference or re-weighting task losses dynamically (Fernando et al., 2023; Lin et al., 2024; Ma et al., 2023). However, these strategies have been predominantly studied in model-free RL or supervised settings, largely outside the domain of latent space planning, where the additional challenge of disentangling dynamics prediction is paramount. Few works have investigated the unique failure modes of multitask learning in this context. We address this gap by analyzing plasticity-related metrics to identify the root cause of performance collapse—an architectural bottleneck in the world model—and propose a novel design to resolve it.
Sparse and Parameter-Efficient Architectures. To overcome the limitations of dense, monolithic models, we turn to sparse and parameter-efficient architectures. Sparse models like Mixture-of-Experts (MoE) (Dai et al., 2024) increase model capacity at a near-constant computational cost through conditional computation. This provides a natural architectural prior for multitask specialization, as different tasks can be routed to different "expert" sub-networks (Obando-Ceron et al., 2024). Concurrently, parameter-efficient fine-tuning (PEFT) methods, notably Low-Rank Adaptation (LoRA) (Hu et al., 2022), enable lightweight, task-specific adjustments. By freezing the bulk of the model and training only a small subset of injected parameters, LoRA can efficiently adapt a model to new tasks or data distributions (Wang et al., 2023; Yang et al., 2025; Zhang et al., 2025). We identify a critical synergy between these two paradigms: MoE provides coarse-grained, input-dependent specialization at an architectural level, while LoRA offers fine-grained, efficient adaptation for targeted plasticity. The integration of these two approaches is a promising but underexplored frontier for building highly adaptable models (Dou et al., 2023; Huang et al., 2024). Our work bridges this gap by unifying MoE and LoRA within a transformer-based world model, creating a system that is both architecturally specialized and dynamically adaptable for large-scale MTRL.
Appendix G Detailed Conclusion and Future Work
G.1 Conclusion
In this paper, we diagnosed the plasticity collapse that plagues unified world models in the context of heterogeneous multi-task planning. Through a principled exploration across the input, model, and optimization dimensions, we introduced ScaleZero, a novel architecture that integrates a sparse Mixture-of-Experts (MoE) backbone. This design creates specialized computational pathways, directly counteracting the representational bottlenecks that lead to performance collapse. Extensive experiments on a diverse suite of 48 tasks across 26 Atari games, 18 DMControl environments, and 4 Jericho worlds show that a single ScaleZero agent matches or surpasses the performance of specialized single-task agents trained from scratch. We also proposed a LoRA-based Dynamic Parameter Scaling (DPS) strategy, which dynamically allocates model capacity in response to task progress. In a proof-of-concept evaluation on DMControl, this strategy achieved competitive performance while reducing total environment interactions by around 28.5%, demonstrating its strong potential for improving sample efficiency in large-scale MTRL.
G.2 Future Work
Our work points to several promising research directions for building more capable and efficient generalist agents:
  • Comprehensive DPS Evaluation: The Dynamic Parameter Scaling (DPS) strategy showed promising results on DMControl. A crucial next step is to rigorously evaluate its performance on more complex and diverse domains like Atari and Jericho. This would involve assessing its generalization capabilities, scalability to a larger number of tasks, and the computational overhead of its adaptive mechanism.
  • Deeper MoE-LoRA Synergy: Our current approach uses MoE for architectural specialization and LoRA for dynamic capacity scaling. A deeper investigation into the synergy between these two could yield further benefits. For instance, one could explore using LoRA to adapt the MoE gating network itself, or using MoE routing decisions to inform which LoRA adapters should be activated, creating a more tightly integrated and adaptive system.
  • Hybrid Online-Offline Learning: Our framework currently operates in a purely online learning setting. A powerful extension would be to combine it with large-scale offline pre-training. By initializing our ScaleZero model with weights from a foundation model pre-trained on vast datasets, we could potentially leverage broad prior knowledge for superior initialization, faster adaptation to new tasks, and higher peak performance, paving the way for truly generalist agents.
We believe these directions offer a clear path toward developing the next generation of multi-task decision-making systems.
Appendix H The Use of Large Language Models
In the preparation of this manuscript, the large language model Gemini 2.5 Pro was employed to assist with language refinement. The primary goal was to enhance the clarity, precision, and readability of the text. The authors have reviewed and approved the final version.